### v1

# references.bib

Abstract
This paper puts forth an innovative approach that fuses deep learning, fractal analysis, and turbulence feature extraction techniques to create abstract artworks in the style of Pollock. The content and style characteristics of the image are extracted by the MindSpore deep learning framework and a pre-trained VGG19 model. An optimisation process is then employed to The method generates high-quality Pollock-style images by combining content loss, style loss and full variance loss to achieve accurate style migration. Furthermore, this paper implements a fractal dimension calculation method based on the difference box-counting method, which effectively estimates the fractal dimension of an image through edge extraction and fractal analysis. The method is based on a two-dimensional discrete wavelet transform using a Haar wavelet to decompose the image in order to extract different frequency information. This is followed by the combination of multiple features to generate unique non-homogeneous token (NFT) labels for the authentication and protection of digital artwork. The experimental results demonstrate that the generated artworks exhibit The method demonstrates significant diversity and complexity in terms of fractal dimensions and turbulence features, while the generated NFT tags ensure the uniqueness and tamperability of each digital collection. The present method organically combines computer vision, digital signal processing and blockchain technology to provide a new solution for the creation and authentication of digital artworks.
Keywords Neural Style Migration, Fractal Analysis, Turbulence Features, NFT Labelling, Deep Learning
1 Introduction
In the domain of contemporary art, abstract art occupies a significant position, characterised by its distinctive mode of expression and profound ideological implications. Jackson Pollock’s Drip Painting has attained the status of a classic within the field of Abstract Expressionism, due to the intricate texture and dynamic expression that characterise it ([ref1, ref25, ref26]). Pollock’s works have a strong visual impact and also present complex fractal characteristics, which provide a basis for digital simulation and scientific analysis of artistic styles.
The advent of artificial intelligence has ushered in a new era of artistic creation, with the application of sophisticated deep learning models such as Generative Adversarial Networks (GANs) and VGG19. These technologies facilitate the simulation of the artistic creation process, enabling the generation of diverse styles and expanding the boundaries of artistic expression. They also provide artists with innovative tools and inspiration. Concurrently, the advent of blockchain technology, particularly the utilisation of non-homogeneous tokens (NFTs) ([ref2, ref3, ref4, ref5]), has established a secure, transparent, and decentralised platform for the authentication and transaction of digital artworks, thereby facilitating the growth of the digital art market.
The objective of this paper is to investigate the potential of combining artificial intelligence, fractal theory and turbulence feature extraction in order to generate abstract artworks in the style of Pollock and achieve unique authentication of digital collections through the use of NFT technology. Specifically, we propose an integrated approach that combines deep learning models, fractal dimension computation and turbulence feature analysis to accurately simulate Pollock’s art style and generate tamper-proof NFT tags via blockchain to ensure the uniqueness of each digital artwork and the verifiability of ownership. This research not only provides a technical path for the creation of digital artworks, but also provides a reliable solution for their authentication and transaction on the blockchain platform, which has both theoretical value and practical application potential.
Figure 1: Pollock Drop Painting
2 Related Work
2.1 Fractal Features of Pollock’s Paintings
In 1999, Richard Taylor’s team made the initial discovery that Pollock’s drip paintings exhibited notable fractal characteristics ([ref6, ref7]). The fractal analysis of Pollock’s works, conducted through the box counting method, yielded fractal dimensions indicative of a self-similar fractal structure, analogous to that observed in natural phenomena. Taylor’s study posited that Pollock may have unconsciously emulated the fractal pattern observed in nature during the creative process. This discovery constituted a pivotal foundation for the subsequent numerical simulation study.
2.2 Application of AI in Art Creation
With the rapid development of AI technology, the application of deep learning models in art creation has attracted increasing attention, particularly the convolutional neural network (CNN)-driven neural style migration technique ([ref8, ref9, ref10]). The neural style migration technique enables the generation of images with different artistic styles by extracting the features of content images and style images. The VGG19 model, a pre-trained deep convolutional network, is a popular choice for advanced feature extraction. In 2016, Leon Gatys et al. proposed a neural style migration method based on the VGG19, which generates a high-quality style through the optimisation of the content loss and the style loss migration images ([ref11, ref12]). This method outperforms Generative Adversarial Networks (GANs) in controlling the generated details and maintaining content consistency, and thus becomes an ideal tool for artistic image processing and is widely used in art creation and digital style simulation.
2.3 Physical Properties in Works of Art
In 2024, Yinxiang et al. demonstrated that the vortex flow characteristics observed in Van Gogh’s Night of the Stars and Moon are highly compatible with the phenomenon of turbulence in physics, particularly in accordance with Kolmogorov’s theory of turbulence ([ref13]). In accordance with the theory, energy is transferred from large-scale vortices to small-scale vortices in a gradual manner, and ultimately dissipated. It was observed that the vortices present in the work not only align with the Kolmogorov turbulence scalar law, but also with the scalar turbulence power spectrum proposed by George Batchelor in 1959.
2.4 NFT Technique
The advent of NFT technology commenced in 2017 ([ref14]), when Dieter Shirley’s team initiated the CryptoKitties project on the Ether blockchain, which facilitated the inaugural large-scale NFT deployment. The FT application ([ref15]) provides a technological guarantee for the uniqueness and ownership authentication of digital assets based on the tamper-proof nature of the blockchain. In 2021, Beeple’s NFT artwork Everydays: The First 5000 Days achieved a high price at Christie’s auction house, indicating that NFT has become a significant presence in the art market. However, the effective integration of the distinctive attributes of artworks with NFT technology remains a topic of ongoing research interest ([ref16]).
3 Methodology
3.1 Algorithm Design
This study proposes a neural style migration algorithm based on the MindSpore deep learning framework ([ref17, ref18]). The algorithm combines the VGG19 model with a variety of loss functions to generate abstract artworks in the Pollock style. The VGG19 model is initially loaded, and its frozen convolutional layer is employed as a feature extractor to obtain multi-layer feature representations from content and style images. With regard to loss function design, the model combines content loss, style loss, total variation (TV) loss, texture loss, and droplet effect loss to generate images that are both artistically styled and naturally smooth.
Content Loss
Content loss is calculated by the mean squared error (MSE) between the generated image and the content image at specified convolutional layer features, ensuring structural consistency. It is defined as:
Content $\text{Content Loss}=\sum_{i}\text{MSE}(C_{i},C_{\text{target}_{i}})$ (1)
where $C_{i}$ represents the features of the generated image at the $i$-th layer, and $C_{\text{target}_{i}}$ denotes the features of the content image at the same layer.
Style Loss
Style loss is based on the Gram matrices of multiple convolutional layers, ensuring the capture of texture and stylistic features of the style image. It is defined as:
Style $\text{Style Loss}=\sum_{j}\text{MSE}(\text{Gram}(S_{j}),\text{Gram}(S_{\text{target}_{j}}))$ (2)
Here, $S_{j}$ represents the features of the generated image at the $j$-th layer, $\text{Gram}(S_{j})$ is its corresponding Gram matrix, and $S_{\text{target}_{j}}$ is the features of the style image at that layer.
Total Variation Loss
Total variation loss is used to suppress noise and smooth the generated image. It is defined as:
TV $\text{TV Loss}=\frac{1}{BCHW}\left(\sum_{b}\sum_{c}\sum_{h}|I_{b,c,h,w}-I_{b,c,h,w+1}|+\sum_{b}\sum_{c}\sum_{w}|I_{b,c,h,w}-I_{b,c,h+1,w}|\right)$ (3)
where $B$ is the batch size, $C$ is the number of channels, and $H$ and $W$ represent the height and width of the image, respectively.
Texture Loss
Texture loss aims to enhance the local texture representation of abstract styles. It is defined as:
Texture $\text{Texture Loss}=\frac{1}{CHW}\left(\text{mean}(\|dx\|^{2})+\text{mean}(\|dy\|^{2})\right)$ (4)
where $dx$ and $dy$ represent the gradients of the feature map in the horizontal and vertical directions, respectively.
Drip Effect Loss
Drip effect loss simulates the ink dripping effect found in Pollock’s art. It is defined as:
Drip Effect $\text{Drip Effect Loss}=-\frac{1}{CHW}\text{mean}(|dy|)+\frac{1}{CHW}\text{mean}(|dy-\text{mean}(dy)|)$ (5)
In this context, $dy$ represents the gradient of the generated image in the vertical direction.
3.2 Fractal Dimension
This study employs the differential box-counting method to quantify the fractal characteristics of the generated artworks, as outlined in references ([ref19, ref20, ref21]). The Sobel filter is employed to extract edge features and perform box counting at varying scales. The number of non-empty boxes in the image is counted, and the box sizes, along with the counting results, are recorded. Subsequently, a linear regression fit is performed on the logarithmic relationship between box size and count value, thereby obtaining the fractal dimension $D$. The fractal dimension $D$ indicates the structural complexity of the image, with larger values signifying more pronounced fractal features.
$D=-\frac{\Delta\log(s)}{\Delta\log(N)}$ (6)
where $N$ is the count of blocks at a specific box size, and $s$ is the box size.
3.3 Wavelet Transform Turbulence Features
To further analyze the turbulence features in the generated artworks, this study employs the two-dimensional discrete wavelet transform (DWT) ([ref22, ref23]), selecting the Haar wavelet as the basis function. This approach captures the local frequency characteristics of the image at different scales and orientations, making it suitable for analyzing turbulence structures and detail variations. The mathematical expression is as follows:
$I(x,y)=\sum_{m}\sum_{n}\psi_{j,k}(x,y)\cdot I(m,n)$ (7)
where $\psi_{j,k}(x,y)$ represents the wavelet basis function, $j$ indicates the scale, and $k$ denotes the position. Through this decomposition, the Haar wavelet can efficiently detect local changes and multi-scale characteristics of turbulence structures within the image.
3.4 Generation of NFT labels
3.4.1 Feature Hashing
In order to guarantee the uniqueness and integrity of the generated digital artefacts, this study employs a methodology that combines multiple fractal spectra with turbulence features to create unique identifiers, designated as Token IDs. In particular, the fractal dimensions and turbulence features are initially integrated into a unified feature vector, which serves as the fundamental descriptive information pertaining to the artwork. Subsequently, the aforementioned feature vector is hashed using cryptographic hashing algorithms, such as SHA-256 ([ref24]), in order to generate a unique Token ID. This not only encodes the distinctive features of the artwork in question, but also provides a means of verifying the artwork’s uniqueness on the blockchain, thereby preventing any instances of forgery or tampering. This mechanism serves to guarantee the authenticity and transparency of digital artworks, whilst simultaneously supporting their traceability and trading in a decentralised network. This, in turn, serves to enhance the trust and value of artworks in the NFT market.
3.4.2 NFT Metadata Generation
The generated Token IDs and their corresponding feature data are integrated into the NFT metadata, thereby ensuring authentication and traceability of digital artworks. The metadata is structured as follows:
  • The Token ID is a unique identifier generated based on a feature hash, which ensures the uniqueness and non-repeatability of each digital collection.
  • Fractal Dimension: an index used to quantify the self-similarity of an image, reflecting its fractal characteristics and highlighting the details and layers of the artwork.
  • Turbulence Characteristics: statistical metrics used to describe the dynamic complexity of an image, capturing its fluidity and complex structure.
  • Generation timestamps: used to record the specific time when the NFT was generated, ensuring its temporal uniqueness and historical traceability.
The metadata is structured in accordance with the NFT standards, such as ERC-721 or ERC-1155, and is deployed on blockchain platforms, including EtherChannel, through the use of smart contracts. The role of smart contracts is to oversee the creation, transfer and validation process for NFTs, thereby ensuring that each NFT is maintained in strict accordance with its associated metadata. The integrity and authenticity of the metadata is assured by the transparency and non-tamperability of the blockchain. Furthermore, the metadata can be stored on-chain or hosted in a decentralised storage solution (e.g. IPFS) to improve data accessibility and persistence. This design not only supports rapid verification and traceability, but also enhances the market value and trust of NFT artefacts.
4 Experimental Results and Analyses
4.1 Style Migration Results
Figure 2: MindSpore-Generated Pollock-Style Paintings
The Pollock-inspired paintings generated on the basis of the MindSpore framework exhibit a multitude of layers of colour and intricate textural structures, effectively mimicking the distinctive style of abstraction evident in Jackson Pollock’s drip paintings. The resulting images evince a clear sense of dynamism and spontaneity, presenting a kaleidoscope of colour tones and interwoven lines. These images fully reflect the fractal and turbulent characteristics of Pollock’s work, thereby proving the high effectiveness of the algorithm in capturing Pollock’s artistic expression techniques.
4.2 Fractal Dimension Results
The fractal dimension, calculated using the difference box counting method, is approximately $D=1.88$, which demonstrates the self-similarity and complex fractal structure of the generated artworks. The estimate of the fractal dimension is highly consistent with the results of Richard Taylor et al. (1999) on Pollock’s paintings, which validates the effectiveness of this study in Furthermore, the fitted curves demonstrate a distinct linear trend, indicating that the box counts of the generated images adhere to a power law distribution, which provides additional support for their fractal characterisation. The image enhanced by the Sobel filter also exhibits a high degree of detail and complex texture structures in the multiscale analysis, thereby substantiating the capacity of the generative algorithm in capturing the intricacies of the drop painting.
Figure 3: Fractal Dimension
4.3 Turbulence Characterisation Results
Figure 4: Turbulence Characterisation Results
The characteristics of turbulence are revealed through the utilisation of a range of power and spectral metrics. The mean value of the turbulence power spectrum extracted using the Haar wavelet transform is $2067.82$, with a variance of $3552.45$. These metrics indicate the overall level of turbulence intensity and characterise the spectral properties and dynamic complexity of the generated images at different scales. The results demonstrate that the generated artworks exhibit significant diversity in spectral distribution, reflecting the complex nature of turbulence. These spectral characteristics complement the analysis of fractal dimension and provide a unique perspective for understanding the dynamic behaviour of the images. Compared to a traditional Fourier transform (FFT), a wavelet transform shows greater sensitivity and accuracy in capturing local frequency features, allowing for a more comprehensive depiction of the complex dynamic behaviour of the images.
4.4 Uniqueness of NFT Tags
The Token ID generated by feature hashing serves to guarantee the uniqueness and unforgeability of each digital collection. Table 1 provides examples of NFT metadata that validate the efficacy of our method in generating reliable and unique NFT labels.
Table 1: Example of NFT metadata
Field Value
token_id a1b2c3d4e5f67890123456789
fractal_dimension 1.88
turbulence_mean_power 2067.82
turbulence_variance_power 3552.45
timestamp 2024-10-27 T12:34:56Z
artist MindSpore-VGG-Pollock
This design guarantees the singularity of each NFT on the blockchain and enables the traceability of its genesis and attributes through metadata. The findings demonstrate that the algorithm employed in this study is highly reliable in generating distinctive and verifiable NFT tags.
5 Conclusion and Discussion
This study proposes a method combining deep learning, fractal analysis, turbulence feature extraction and blockchain NFT technology for the generation of digital artworks in the style of Pollock. The method in question achieves an accurate simulation of the style and structure of the artwork in question by performing style migration through the VGG19 neural network, calculating the fractal dimension based on the difference box-counting method, and extracting turbulence features using the wavelet transform. The experiments demonstrate that the generated artworks exhibit the defining characteristics of Pollock’s style, including fractal structure and turbulence characteristics. Additionally, the generated NFT tags ensure the uniqueness and tamperability of the digital artworks, providing a novel technical solution for the digital art market.
The innovation of this research lies in the combination of artificial intelligence and blockchain technology, which serves to broaden the expression and authentication of digital art. The distinctive tag generated by NFT technology guarantees the singularity and proprietorship traceability of each digital artwork on the blockchain, thereby ensuring the security of the circulation and protection of digital artworks. Furthermore, the thesis corroborates the efficacy of fractal analysis and turbulence feature extraction techniques in digital art creation, thereby enhancing the expressive potential of modern artworks.
Further research could extend the generation and analysis methods for different art styles, with a particular focus on the digital simulation of other classical styles. Furthermore, the generation and transaction processes associated with NFT labels can be optimized to enhance efficiency and market activity, thereby facilitating the growth of the digital art market. Furthermore, by fostering interdisciplinary collaboration between technology and the arts, and delving deeply into the potential applications of technology in artistic creation, a wealth of opportunities will be unlocked, propelling the innovative evolution of digital art.
Acknowledgment
Thanks for the support provided by MindSpore Community.
Declaration of Code and Tool Availability:
Our results can be obtained through experiments using the MindSpore framework. The code for this article is available at https://github.com/mindspore-lab/models/tree/master/research/arxiv_papers.
[…]